Thursday, December 8, 2011

Muddiest Point!!!

Hey guys its Shane, and this is my muddiest point. The thing I'm confused on is the difference between the genotypic ratio and the parents genotypes. I'm kind of counfused on which one to put down when it says parents genotypes and if you could help that would be cool.

Wednesday, December 7, 2011

Me Es Scribe!

Today In our shortened class, after we checked in the homework on genetic disorder problems, we took a quiz on 14.2 which involved  disorders caused by individual genes which include Sickle Cell Disease, Cystic Fibrosis, and Huntington's Disease. The quiz also touched on Nondisjunction- which is an error that occurs during meiosis where the homologous chromosomes fail to separate properly.

After the quiz we went over the diseases that were caused by Autosomal dominant alleles (Huntington's disease) and Co- Dominant alleles.

At the end of class we worked with the laptops on a web quest about karyotyping and chromosomal abnormalities.

The next scribe is Jonathan

Tuesday, December 6, 2011

Scribe o da day

Hey guys it's Shane, and in class today we learned about AUTOSOMAL RECESSIVE DISORDERS and PROBABILITY!!!!

Probability consists of three different things. One of them is EXPECTED. This means the fifty fifty chance. Say you were to flip a coin ten times, you would expect it to land on heads five times and it to land on tails five times. But if you were to ACTUALLY flip the coin it more likely not land five and five. The difference between expected and actual probability would be called the DEVIATION. Deviation also means that if you have more tries than the percent will more likely be close to fifty percent for each side. Say if you were to have had three sons the probability would be 1/8th.

1/2 * 1/2 * 1/2=1/8

You would multiplie 1/2*1/2*1/2 because it is a 1/2 chance of getting a son. As you can see that the answer was smaller because of DEVIATION.

Autosomal Recessive Disorders are NON-SEX disorders, since AUTOSOMAL means non-sex linked. When you have a autosomal recessive disorder you have to have two recessive alleles to be able to have the disorder. BUT, for Autosomal DOMINANT disorders you can have either one or two dominant alleles to have the disease.




The next scribe is Michael M...

Monday, December 5, 2011

What's the Homework? Oh, more punnet square problems...

Hey what's up it's Yasmin! I was the scribe for FRIDAY. So We learned how to solve problems with multiple alleles. It wasn't easy because we were used too solving problems that only had two possible alleles for an answer. Like i said before,It wasn't EASY but we learned and did our best to complete the problems. Nothing is impossible but while i was doing some of these problems i thought they were impossible, but that's just me. Furthermore, we also worked on Sex linkage problems. To keep it short and simple(i don't want to bore you) a Sex-link trait is where a gene is carried ONLY on the X chromosome. The picture is an example of a multiple allele problem.






Pedigrees!!




Today in class, we learned about PEDIGREES!

Pedigrees are used to trace inheritance through families. So basically it's what people use to find out who had what traits in a family, and what probability other people in the family have of getting that trait.


The picture below is a pedigree for the inheritance of red-green color blindness. The boxes are boys, and the circles are girls. Also the fully shaded boxes/circles mean that the person is colorblind. While the half shaded ones mean that they carry the trait for it, but it is not dominant. The empty ones mean that they don't have the trait at all


next scribe is: Shane

Thursday, December 1, 2011

ʇsod ǝqıɹɔs s,ʎɐpsɹnɥʇ


WHOAAAA :O I don't know if you noticed but the title is "thursday's scribe post" UPSIDE DOWN! Cool, I know. So, Anka here obviously! Here's what we did in class today, Thursday, the FIRST DAY OF DECEMBER (see what I did there, Christmas AND Hanukkah colors)! WOOTWOOT!
We started off class by correcting our homework, UP 29-31, where we learned how to efficiently cross involving one trait (monohybrid crosses). You had to make sure you knew the difference between homozygous and heterozygous traits. If it asked for homozygous green pods, you had to write, GG, as opposed to heterozygous green pods, Gg, and recessive yellow pods, gg. Then you had to seperate those homologous pairs which results in the seperation of alleles. This process is called segregation of meiosis, as shown in the picture. After that you formed and completed a punnett square and moved on to determine the genotypic and phenotypic ratios. Afterwards, our teacher gave us another sheet with a monohybrid cross to be completed to ensure we could all do it!
Then we went on to watch and take notes on a powerpoint. We talked about Independent Assortment (Mendels law), the process of which alleles seperate independently from eachother which then leads to the formation of gametes. Next, we talked about Incomplete Dominance which is when a cross between two phenotypes produces a third phenotype that is considered a BLENDING of the two different traits.
Ex. straight x curly = wavy
Finally, we talked about Co-Dominance which is when both alleles are dominant and expressed equally (NOT BLENDED).
Ex. white rose x red rose = pink rose
The very last thing we did in class was complete an example of Incomplete Dominance together as a class and start on our homework, UP 33-41.
GOOD LUCK ON YOUR HOMEWORK, it's not bad at all!

NEXT SCRIBE IS.............. YASMIN(: [SORRY] gooooodluck, gurrrrl.

Scribe Of The Day

Hola Clase!
in class on wednesday, we went to el lab de lenguas and took a test there. we had to talk into the microphone about a pictures that we saw, and say what the people in it were doing. We also had a sheet for homework where we had to label body parts on a person.

next scribe is: Jordyn